OpenStudy (anonymous):

Which set of numbers represents a Pythagorean triple? a; 6, 9, 12 b; 7, 10, 12 c; 16, 18, 25 d; 27, 36, 45

OpenStudy (anonymous):

can someone explain to me how to do this?

OpenStudy (anonymous):

A Pythagorean Triple is a set of 3 numbers that can represent the three sides of a RIGHT triangle. For example, 3,4,5 can represent the sides of a right triangle becuase 3^2 + 4^2 = 5^2 9 + 16 = 25 SO go through the choices and determine which set of numbers will a^2 + b^2 = c^2.